OpenStudy (anonymous):

If cos theta = 1/3 and 0 is less than or equal to theta, which is less than or equal to pi/2, find tan theta.

OpenStudy (anonymous):

\[If \cos \theta = 1/3 and 0\le \theta \le \pi/2, find \tan \theta. \]